OpenStudy (anonymous):

simplify: x+1/ x^2-1

jimthompson5910 (jim_thompson5910):

x^2 - 1 factors to ????

OpenStudy (anonymous):

(x-1) (x+1)

jimthompson5910 (jim_thompson5910):

This makes \[\Large \frac{x+1}{x^2-1}\] turn into \[\Large \frac{x+1}{(x-1)(x+1)}\] and the "x+1" terms will cancel out leaving you with \[\Large \frac{1}{x-1}\]

OpenStudy (anonymous):

yay thanks again!!

jimthompson5910 (jim_thompson5910):

sure thing